# Intuitive Surgical SurgToolLoc Challenge Results: 2022-2023

### v3

@@ Abstract
Robotic assisted (RA) surgery promises to transform surgical intervention. Intuitive Surgical is committed to fostering these changes and the machine learning models and algorithms that will enable them. With these goals in mind we have invited the surgical data science community to participate in a yearly competition hosted through the Medical Imaging Computing and Computer Assisted Interventions (MICCAI) conference. With varying changes from year to year, we have challenged the community to solve difficult machine learning problems in the context of advanced RA applications. Here we document the results of these challenges, focusing on surgical tool localization (SurgToolLoc). The publicly released dataset that accompanies these challenges is detailed in a separate paper https://arxiv.org/abs/2501.09209 [1].
@@ 1 Introduction
Robotic-assisted (RA) surgeries become more prevalent with each passing year [2, 3]. Importantly, RA surgeries also generate rich streams of data, e.g. video recordings, tool kinematics, and events gathered from the robotic system, that lend themselves to quantitative analyses. Pursuits such as the quantification of surgical performance, autonomous analysis of procedures, surgical skill evaluation, AI-guided surgical planning, and surgical data science, in general, can all exploit this rich source of data [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. However, despite the ability to collect large amounts of it, annotating it to provide meaningful context remains a major bottleneck in the field.
@@ 3.1.6 Submission Process
Each team developed its own algorithm following the guidelines provided by the GC platform [^4] https://grand-challenge.org/documentation/algorithms/ and the organizing team of the challenge. Every submitted algorithm was configured to receive as input either a preliminary test dataset containing two videos for debugging purposes or the final test dataset described in 4.1.3. The algorithm’s output had to follow a standardized file format for each category. For category 1 -surgical tool classification-, the output file had to generate a list of dictionaries containing information on the tools present in each frame of the input video in a boolean format. For category 2 - Surgical tool classification and localization-, the algorithm had to output a dictionary containing the set of tools detected in each frame with its corresponding bounding box cartesian coordinates (x, y). Details on the creation of the algorithms, samples of input data, and the required format for output files were provided to the participating teams  [^5] https://surgtoolloc.grand-challenge.org/submission/.
@@ 11 ITeM Furtwangen University, University of Lipzig Germany 1 & 2 Y
12 MM Chinese University of Hong Kong, National University of Singapore, University College London Hong Kong, Singapore, UK 1 only Y
@@ 3.10.1 Method Description
For both tasks, the starting point was to extract feature vectors exploiting a self-supervised learning approach based on DINO (cf. figure 15 for a schematic illustration). Prior to training, frames were resized to $308\times 208$ px and frame intensities normalized to ($-1,1$). A DINO model, pre-trained on the ImageNet data, was trained on the given pre-processed training data with slightly adapted default DINO-parameters.
@@
A temporal global crop was applied to consider the temporal components of the input data. This approach involves randomly selecting a global crop from a frame that is in temporal correspondence with the input frame, with a tolerance of ±$\pm 5$ s. This approach ensures temporal consistency between the input frame and the selected global crop, i.e. the temporal structure of the data is preserved which further diversifies the input data.
@@ 3.11.1 Method Description
      $PER=\frac{\sum_{i}^{h}\sum_{i}^{w}x_{i,j}}{h.w}\end{aligned}\qquad\Longleftrightarrow\qquad\begin{aligned} x_{i,j}=1&\text{ for }x>0.9\\ x_{i,j}=0,&\text{ Otherwise}\end{aligned}$
      $NER=\frac{\sum_{i}^{h}\sum_{i}^{w}x_{i,j}}{h.w}\end{aligned}\qquad\Longleftrightarrow\qquad\begin{aligned} x_{i,j}=1&\text{ for }x<0.1\\ x_{i,j}=0,&\text{ Otherwise}\end{aligned}$
@@ 3.12.1 Method Description
Figure 18: MM: R50+ViT-B_16 hybrid model
@@ 3.12.2 Model Training
Figure 19: MM: The training process.
@@ 3.14.2 Implementation Details
During the training phase of the network, the video frames were resized to a height of 270 and a width of 480. Subsequently, the upper-middle regions (shaped in 240$\times$336) of the resized frames were cropped as input to the network. This cropping process removed the black regions present on both sides and the text labels located at the bottom of the frames. Each training sample for one video was formed by splitting it into 32 segments and randomly sampling one frame from each segment. At the network test phase, the continuous regions with pixel values less than 1, which represented the black background, were first identified and removed from the frames. The remained foreground was then resized to 270$\times$336, from which the upper 240$\times$336 regions were cropped as input to the network. All frames of a video were taken as input for making predictions at the test phase. The input RGB frames were normalized using the means of (0.432, 0.395, 0.376) and the standard deviations of (0.228, 0.221, 0.217). The network outputted the existence probabilities of all 14 tools, from which the first 3 classes having probabilities larger than 0.5 were selected as the predicted tool labels.
The framework employed the first four convolutional blocks of a ResNet-18 model pre-trained on ImageNet as the feature extractor. To preserve the high spatial resolution of the output feature maps, the max pooling layer following the first convolutional block is eliminated. The resulting feature maps have a spatial shape of 15$\times$21 and 256 channels for a 336$\times$240 frame. The MLP for post-processing preserves the feature channel number at 256. These features are finally fed through a linear transformation network and a Sigmoid activation layer to predict frame-wise tool existence probabilities.
@@ 4.3.2 Model Training
We again use a FasterRCNN, initialized with the weights from the model pretrained on EndoVis2017. We batch together images from the weak and the strongly labeled datasets, employing class balanced sampling to ensure classes are more evenly represented. We apply random photo-metric and spatial augmentations to all the images. The strong supervision loss is calculated as the standard FasterRCNN loss summed and the weak supervision loss calculated using binary cross entropy. The model parameters were optimised using stochastic gradient descent with learning rate of $1$$\times$$10^{-3}$ , moment $0.9$ and weight decay of $1$$\times$$10^{-4}$. We reserved 10% of our pseudo labeled data to use as a validation set for early stopping.
@@ 4.5.2 Model Training
To improve the quality of training data, a morphological opening and closing was applied to the segmentation mask to remove noise and close holes after running TernausNet. The images with poor segmentation masks were removed, via comparing the size of the segmentation blob with certain thresholds. Segmentation blobs with areas $>0.02\times\textrm{total image area}$ were kept and segmentation masks with more than 40% predicted foreground pixels in the image were removed. Furthermore, the team also down-sampled given images to the size of $512\times 512$ with 0 mean and unit standard deviation. The train-test split was 70-30. The initial learning rate was set to 1 changing at every epoch with cosine annealing and the optimizer was stochastic gradient descent (SGD). The loss function was binary cross entropy. An NVIDIA RTX4090 GPU was used for the model training.
@@ 4.7.2 Method
After generating the preliminary round pseudo-labeled datasets, we trained a detection model called $Det_{tools}$ using YOLOv8 [85]. Similarly, we performed inference on the sampling dataset using the trained $Det_{tools}$ model to obtain detection boxes $BBox_{tools}$ corresponding to instruments in each video frame (localizing all instruments in every frame). Since both $Det_{parts}$ and $Det_{tools}$ contain positional information of instruments in the video frames, but the first round of pseudo-labels had considerable noise, $Det_{tools}$ had lower accuracy. Therefore, we adopted a multi-round label filtering approach to select high-quality pseudo-labels, as shown in Algorithm 2 (Fig. 42). For each detection box $bbox_{tool}$ from $BBox_{tools}$ that corresponds to a particular instrument category, we checked if there exist any $bbox_{part}$ with category tip that has IoU $>$ 0.8 with $bbox_{tool}$, and any $bbox_{part}$ with category clevis that has IoU $>$ 0.8 with $bbox_{tool}$. Only when all $bbox_{tool}$ in a frame satisfy these conditions, the pseudo-label is considered trustworthy.

### v4

# Intuitive Surgical SurgToolLoc and SurgVU Challenges Results: 2022-2025

@@ Abstract
Robotic assisted (RA) surgery promises to transform surgical intervention. Intuitive Surgical is committed to fostering these changes and the machine learning models and algorithms that will enable them. With these goals in mind we have invited the surgical data science community to participate in a yearly competition hosted through the Medical Imaging Computing and Computer Assisted Interventions (MICCAI) conference. With varying changes from year to year, we have challenged the community to solve difficult machine learning problems in the context of advanced RA applications. Here we document the results of these challenges, focusing on surgical tool localization (SurgToolLoc) and surgical visual understanding (SurgVU). The publicly released dataset that accompanies these challenges is detailed in a separate paper https://arxiv.org/abs/2501.09209 [1].
@@ 1 Introduction
Robotic-assisted (RA) surgeries become more prevalent with each passing year [2, 3]. Importantly, RA surgeries also generate rich streams of data, e.g. video recordings, tool kinematics, and events gathered from the robotic system, that lend themselves to quantitative analyses. Pursuits such as the quantification of surgical performance, autonomous analysis of procedures, surgical skill evaluation, AI-guided surgical planning, and surgical data science, in general, can all exploit this rich source of data [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. However, despite the ability to collect large amounts of it, annotating it to provide meaningful context remains a major bottleneck in the field.
@@ 3.1.6 Submission Process
Each team developed its own algorithm following the guidelines provided by the GC platform [^4] https://grand-challenge.org/documentation/algorithms/ and the organizing team of the challenge. Every submitted algorithm was configured to receive as input either a preliminary test dataset containing two videos for debugging purposes or the final test dataset described in 3.1.5. The algorithm’s output had to follow a standardized file format for each category. For category 1 -surgical tool classification-, the output file had to generate a list of dictionaries containing information on the tools present in each frame of the input video in a boolean format. For category 2 - Surgical tool classification and localization-, the algorithm had to output a dictionary containing the set of tools detected in each frame with its corresponding bounding box cartesian coordinates (x, y). Details on the creation of the algorithms, samples of input data, and the required format for output files were provided to the participating teams  [^5] https://surgtoolloc.grand-challenge.org/submission/.
@@ 11 ITeM Furtwangen University, University of Lipzig Germany 1 & 2 Y
12 $M$ Chinese University of Hong Kong, National University of Singapore, University College London Hong Kong, Singapore, UK 1 only Y
@@ 3.10.1 Method Description
For both tasks, the starting point was to extract feature vectors exploiting a self-supervised learning approach based on DINO (cf. figure 15 for a schematic illustration). Prior to training, frames were resized to $308\times 208$ px and frame intensities normalized to (1,1-1,1). A DINO model, pre-trained on the ImageNet data, was trained on the given pre-processed training data with slightly adapted default DINO-parameters.
@@
A temporal global crop was applied to consider the temporal components of the input data. This approach involves randomly selecting a global crop from a frame that is in temporal correspondence with the input frame, with a tolerance of $\pm 5$ s. This approach ensures temporal consistency between the input frame and the selected global crop, i.e. the temporal structure of the data is preserved which further diversifies the input data.
@@ 3.11.1 Method Description
      $\begin{aligned} PER=\frac{\sum_{i}^{h}\sum_{i}^{w}x_{i,j}}{h.w}\end{aligned}\qquad\Longleftrightarrow\qquad\begin{aligned} x_{i,j}=1&\text{ for }x>0.9\\ x_{i,j}=0,&\text{ Otherwise}\end{aligned}$
      $\begin{aligned} NER=\frac{\sum_{i}^{h}\sum_{i}^{w}x_{i,j}}{h.w}\end{aligned}\qquad\Longleftrightarrow\qquad\begin{aligned} x_{i,j}=1&\text{ for }x<0.1\\ x_{i,j}=0,&\text{ Otherwise}\end{aligned}$
@@ 3.12.1 Method Description
Figure 18: $M$: R50+ViT-B_16 hybrid model
@@ 3.12.2 Model Training
Figure 19: $M$: The training process.
@@ 3.14.2 Implementation Details
[…]
@@ 4.3.2 Model Training
[…]
@@ 4.5.2 Model Training
To improve the quality of training data, a morphological opening and closing was applied to the segmentation mask to remove noise and close holes after running TernausNet. The images with poor segmentation masks were removed, via comparing the size of the segmentation blob with certain thresholds. Segmentation blobs with areas $0.02\times\textrm{total image area}$ were kept and segmentation masks with more than 40% predicted foreground pixels in the image were removed. Furthermore, the team also down-sampled given images to the size of $512\times 512$ with 0 mean and unit standard deviation. The train-test split was 70-30. The initial learning rate was set to 1 changing at every epoch with cosine annealing and the optimizer was stochastic gradient descent (SGD). The loss function was binary cross entropy. An NVIDIA RTX4090 GPU was used for the model training.
@@ 4.7.2 Method
After generating the preliminary round pseudo-labeled datasets, we trained a detection model called $Det_{tools}$ using YOLOv8 [86]. Similarly, we performed inference on the sampling dataset using the trained $Det_{tools}$ model to obtain detection boxes $BBox_{tools}$ corresponding to instruments in each video frame (localizing all instruments in every frame). Since both $Det_{parts}$ and $Det_{tools}$ contain positional information of instruments in the video frames, but the first round of pseudo-labels had considerable noise, $Det_{tools}$ had lower accuracy. Therefore, we adopted a multi-round label filtering approach to select high-quality pseudo-labels, as shown in Algorithm 2 (Fig. 42). For each detection box $bbox_{tool}$ from $BBox_{tools}$ that corresponds to a particular instrument category, we checked if there exist any $bbox_{part}$ with category tip that has IoU >> 0.8 with $bbox_{tool}$, and any $bbox_{part}$ with category clevis that has IoU >> 0.8 with $bbox_{tool}$. Only when all $bbox_{tool}$ in a frame satisfy these conditions, the pseudo-label is considered trustworthy.